# Addressing divergent representations from causal interventions on neural networks

### v3

@@ 5 How might we avoid divergent representations?
We have thus far shown that divergent representations are empirically common and that they can be harmless or harmful depending on multiple factors. We now consider the question of how such divergences might be avoided. Some existing methods solve this by constraining counterfactual features (in our case, intervened representations) directly to the natural manifold (verma2024counterfactualexplanationreview). However, we seek a method that generates principled interventions that are constrained to be innocuous. In this pursuit, we first apply the Counterfactual Latent (CL) loss from grant2025mas to the DAS experiments from wu2024alpacadas and find that we can reduce intervened divergence while preserving accuracy on a Llama based LLM. We then introduce a modified CL loss that only targets causal dimensions, and we show that it can improve OOD intervention accuracy.
@@ 5.2 Modified CL loss improves OOD interventions in synthetic model settings
We hypothesized that reduced divergence could improve IIA when transferring the DAS rotation matrix to OOD settings. To test this, we partitioned the synthetic task into a dense and a sparse cluster of classes based on the relative spacing of 4 distinct classes within each cluster (see Appendix A.4.4 for partition details). We then performed a separate MLP and DAS training on each partition and evaluated the IIA on the held-out partition to produce Figure 3F. The CL loss performs better than the behavioral loss in these OOD settings. We provide CL loss ablations in Appendix A.4.5.
@@ 6 Discussion and limitations
Where does this leave us with respect to causal interventions in mechanistic interpretability? Given our theoretical findings, any divergence outside of the null-space of NN layers is potentially pernicious. This poses challenges for aspirations of a complete mechanistic understanding of NNs using existing methods alone. However, we note that many practical mechanistic projects can be satisfied by collecting sufficiently large intervention evaluation datasets, and continued development of methods such as the CL loss can reduce the problem even further. We are optimistic for future developments in this field.
@@ 8 LLM Usage Statement
We used ChatGPT to elicit editing suggestions for clarity, as well as to improve notational consistency and clarity in the hidden pathways and dormant behavioral changes formalizations in Sections 4.1 and 4.2. We also used ChatGPT to provide an initial layout of the proof offered in Appendix A.2 showing that axis-aligned hyperrectangles are the only manifold shape that do not have divergent source-target vector pairs in coordinate patching settings.
@@ A.1.2 Measuring Divergence
For each reference point $x_{i}$, we identified its $k$ nearest neighbors $\mathcal{N}_{k}(x_{i})$ in Euclidean space and computed the local tangent subspace via PCA. Let $U_{d}(x_{i})\in\mathbb{R}^{D\times d}$ denote the top $d$ principal components explaining at least 95% of the local variance. The projection matrix onto this tangent space is $P_{i}=U_{d}U_{d}^{\top}$.
@@ A.2.1 Activation patching in balanced subspaces
A.3 CL loss applied to Boundless DAS
@@
A.4 CL Loss in Synthetic Settings
@@
A.4.1 Synthetic Dataset Construction
@@ A.4.3 DAS Training
A.4.4 Out-of-Distribution Experimental Details
@@
A.4.5 Further CL Loss Explorations

### v4

@@ 5 How might we avoid divergent representations?
We have thus far shown that divergent representations are empirically common and that they can be harmless or harmful depending on multiple factors. We now consider the question of how such divergences might be avoided. Some existing methods solve this by projecting counterfactual features (in our case, intervened representations) directly to the natural manifold (verma2024counterfactualexplanationreview). However, we seek a method that generates principled interventions that are constrained to be innocuous. In this pursuit, we first apply the Counterfactual Latent (CL) loss from grant2025mas to the DAS experiments from wu2024alpacadas and find that we can reduce intervened divergence while preserving accuracy on a Llama based LLM. We then introduce a modified CL loss that only targets causal dimensions, and we show that it can improve OOD intervention accuracy.
@@ 5.2 Modified CL loss improves OOD interventions in synthetic model settings
What is the practical utility of reducing divergence? We hypothesized that divergence could influence IIA when transferring the DAS alignment to OOD settings. To test this, we partitioned the synthetic task into a dense and a sparse cluster of classes based on their relative spacing (Appx. A.5.4). We then trained an MLP and alignment on each partition and evaluated the alignment on the held-out partition. The CL loss performed better than the behavioral loss in these OOD settings (Figure 3F). We then regressed OOD IIA on training EMD to find an anti-correlation (coef. -.34, $R^{2}=.73$, F(1,28)=75.28, $p<.001$), showing that divergence can predict lower OOD performance (Appx. A.6).
@@ 6 Discussion and limitations
Where does this leave us with respect to causal interventions in mechanistic interpretability? Given our theoretical findings, any divergence outside of the null-space of NN layers is potentially pernicious. This poses challenges for aspirations of a complete mechanistic understanding of NNs using existing methods alone. However, we note that many practical mechanistic projects can be satisfied by collecting sufficiently large intervention evaluation datasets, and continued development of methods such as the CL loss can reduce the problem even further. We are optimistic for the future of this field.
@@ 8 LLM Usage Statement
We used ChatGPT to generally edit for clarity, as well as to improve notational consistency in the behavioral null-space, hidden pathways, and dormant behavioral changes formalizations in Sections 4.1 and 4.2. We also used ChatGPT to provide an initial layout of the proof offered in Appendix A.2 showing that axis-aligned hyperrectangles are the only manifold shape that do not have divergent source-target vector pairs in coordinate patching settings, and we used it to suggest, implement, and generate the initial writeup for the additional divergence measures in Appendix A.1.2.
@@ A.1.2 Measuring Divergence
For each reference point $x_{i}$, we identified its $k$ nearest neighbors $\mathbb{N}_{k}(x_{i})$ in Euclidean space and computed the local tangent subspace via PCA. Let $U_{d}(x_{i})\in\mathbb{R}^{D\times d}$ denote the top $d$ principal components explaining at least 95% of the local variance. The projection matrix onto this tangent space is $P_{i}=U_{d}U_{d}^{\top}$.
@@ A.2.1 Activation patching in balanced subspaces
A.3 Practical Algorithm for Harmless Divergence when Behavioral Null-Space Characterizes the Full Harmless Set
In settings where perturbations $v\in\mathcal{N}(\psi,X)$ are treated as harmless and perturbations $v\notin\mathcal{N}(\psi,X)$ as harmful, the behavioral null-space formalism suggests a practical procedure for testing the harmlessness of a given divergence $v$. Let $X_{K}\subset\mathbb{R}^{d}$ be the set of natural representations for class $K$, and let $\hat{x}_{K}\in\mathbb{R}^{d}$ be an intervened representation for class $K$. To approximate the natural manifold $\mathcal{M}_{K}$ locally around $\hat{x}_{K}$, we first select the $n$ nearest neighbors of $\hat{x}_{K}$ in $X_{K}$:
$\mathbb{N}_{n}(\hat{x}_{K})=\{x^{(1)},\dots,x^{(n)}\}\subset X_{K}.$
Let $U\in\mathbb{R}^{n\times d}$ be the matrix whose rows are the neighbors $u_{i}^{\top}=(x^{(i)})^{\top}$, and let
$\mu_{K}=\frac{1}{n}\sum_{i=1}^{n}x^{(i)}\in\mathbb{R}^{d}$
denote their mean. Define the centered data matrix
$\tilde{U}=\begin{bmatrix}(x^{(1)}-\mu_{K})^{\top}\\ \vdots\\ (x^{(n)}-\mu_{K})^{\top}\end{bmatrix}\in\mathbb{R}^{n\times d}.$
We compute a rank-$r$ PCA of $\tilde{U}$, obtaining the top $r$ principal directions $Q_{r}\in\mathbb{R}^{d\times r}$ (columns are orthonormal). The corresponding local projection operator is
$\Pi_{K}(x)=\mu_{K}+Q_{r}Q_{r}^{\top}(x-\mu_{K}).$
The local projection of the intervened representation is then
$\hat{x}_{\mathrm{proj}}=\Pi_{K}(\hat{x}_{K}),$
and we define the divergence vector as
$v=\hat{x}_{K}-\hat{x}_{\mathrm{proj}}.$
We now provide Algorithm 1 as a practical method for classifying divergence as harmless or pernicious. We note, however, this algorithm only approximates harmlessness and is not guaranteed to be successful. This is especially the case for situations prone to dormant behavioral changes (Sec. 4.2.2).
Algorithm 1 Classifying the harmlessness of a divergence vector when the behavioral null-space characterizes harmlessness
1:Intervened representation $\hat{x}_{K}\in\mathbb{R}^{d}$ for class $K$; natural class representations $X_{K}\subset\mathbb{R}^{d}$; evaluation set $X_{\mathrm{eval}}\subset\mathbb{R}^{d}$; function $\psi:\mathbb{R}^{d}\to\mathbb{R}^{d^{\prime}}$; neighborhood size $n$; local dimension $r$; tolerance $\epsilon\geq 0$.
2:Classification of the divergence vector $v$ as harmless or harmful.
3:(Local manifold estimation for class $K$) Let $\mathbb{N}_{n}(\hat{x}_{K})=\{x^{(1)},\dots,x^{(n)}\}\subset X_{K}$ be the $n$ nearest neighbors of $\hat{x}_{K}$ in $X_{K}$.
4:Compute the mean $\mu_{K}=\frac{1}{n}\sum_{i=1}^{n}x^{(i)}$.
5:Form the centered matrix $\tilde{U}\in\mathbb{R}^{n\times d}$ with rows $(x^{(i)}-\mu_{K})^{\top}$.
6:Perform rank-$r$ PCA on $\tilde{U}$ to obtain the top $r$ principal directions $Q_{r}\in\mathbb{R}^{d\times r}$.
7:Define the local projection $\Pi_{K}(x)=\mu_{K}+Q_{r}Q_{r}^{\top}(x-\mu_{K})$ and set $\hat{x}_{\mathrm{proj}}\leftarrow\Pi_{K}(\hat{x}_{K})$.
8:Define the divergence vector $v\leftarrow\hat{x}_{K}-\hat{x}_{\mathrm{proj}}$.
9:(Behavioral test over a broader context) For each $x\in X_{\mathrm{eval}}$, compute $\Delta(\psi,x)=\lVert\psi(x+v)-\psi(x)\rVert$.
10:if $\max_{x\in X_{\mathrm{eval}}}\Delta(x)\leq\epsilon$ then
11:  return $v$ is Harmless.
12:else
13:  return $v$ is Harmful.
14:end if
A.4 CL loss applied to Boundless DAS
@@
A.5 CL Loss in Synthetic Settings
@@
A.5.1 Synthetic Dataset Construction
@@ A.5.3 DAS Training
A.5.4 Out-of-Distribution Experimental Details
@@
A.5.5 Further CL Loss Explorations
@@
A.6 Linear Regression
In an effort to establish a more general, concrete relationship between intervened divergence an out-of-distribution (OOD) intervention performance, we performed a linear regression on trained alignments using EMD along causal axes (as discovered through the alignment training) as the independent variable and interchange intervention accuracy (IIA) as the dependent variable. We performed these regressions independently on the Default task and the OOD task trainings, each training consisting of two partitions with  5training seeds and 3 types of alignment trainings: DAS behavioral loss only, CL loss only, and DAS+CL loss, creating  30trainings total. We used the statsmodels python package (seabold2010statsmodels) to perform the regression.
Dep. Variable: IIA R-squared: 0.729
Model: OLS Adj. R-squared: 0.719
Method: Least Squares F-statistic: 75.28
Date: Wed, 26 Nov 2025 Prob (F-statistic): 2.00e-09
Time: 11:54:42 Log-Likelihood: 76.575
No. Observations: 30 AIC: -149.2
Df Residuals: 28 BIC: -146.3
Df Model: 1
Covariance Type: nonrobust
coef std err t P>|>|t|| [0.025 0.975]
Intercept 0.9885 0.004 243.666 0.000 0.980 0.997
Training EMD -0.3424 0.039 -8.677 0.000 -0.423 -0.262
Omnibus: 33.330 Durbin-Watson: 1.903
Prob(Omnibus): 0.000 Jarque-Bera (JB): 98.629
Skew: -2.235 Prob(JB): 3.83e-22
Kurtosis: 10.676 Cond. No. 11.1